Hans is buying groceries. He buys 2 pounds of ground beef, 3 apples, and 2 loaves of bread. A pounds of ground beef cost $3.40. Apples are 3 for $1. Each loaf of bread costs $1.45. The sales tax is 4%. Which of the following represents the total cost of Hans’ groceries?
A) 6.8 (0.04) +1 x (0.04) + 2.9 (0.04)
B) (.4) (6.8+ 1 + 2.9)
C) (0.04) (6.8 + 1 + 2.9)
D) (1.04) (6.8 + 1 + 2.9)
The handbook say the answer D, but I am bit puzzled because I manually solve the addition problem of ground beef, apples, and bread. I don’t understand how sale tax from .04 became “1.04”. That part is confusing me.

Answer by greenestamps(13200)



The expression in parentheses in answer choices B, C, and D is the cost of the groceries before tax.

Answer choice C show the amount of the tax -- 4% of the cost of the groceries: .04(6.8+1+2.9).

The total cost of the groceries is the cost of the groceries themselves plus the tax:
1(6.8+1+2.9) + .04(6.8+1+2.9) = (1+.04)(6.8+1+2.9) = (1.04)(6.8+1+2.9), which is answer D.